# latexText page

### v1

@@ 4.1.5 Applying Grad-ECLIP to explain CNN-based CLIP
Although the methodology in §3.2 for Grad-ECLIP is derived from CLIP with Transformer architecture, here we show that our method is also applicable to CNN-based CLIP by using the attention layer in the final attention pooling. Figure 7 shows the visual explanation results for ResNet50-CLIP using Grad-ECLIP, and compared to other explanation methods that are compatible with CNN-based CLIP, including Grad-CAM, CLIPSurgery, and RISE. From the examples, our Grad-ECLIP is able to generate clear and reasonable text-specific heat maps for interpreting CNN-based CLIP. In contrast to our method, Grad-CAM can produce similar highlight on the explained objects, but its heat map is more noisy in the background. Similar to its performance on ViT-based CLIP, CLIPSurgery generates rougher heat maps, which tend to put high values on a coarse region of the object. Meanwhile, RISE fails to explain CLIP model.
@@
TABLE I: Faithfulness evaluation of image explanation on the ImageNet validation set: AUC for Deletion and Insertion curves, based on Top-1 (@1) or Top-5 (@5) classification accuracy. Either the ground-truth or the prediction are used as the text input into CLIP. The second best is shown with underline.
@@
Ground-truth Prediction Ground-truth Prediction
Method @1 @5 @1 @5 @1 @5 @1 @5
raw attention 0.3831 0.6239 - - 0.2492 0.4195 - -
Rollout 0.4082 0.6556 - - 0.2803 0.4665 - -
Grad-CAM 0.3417 0.5628 0.3518 0.5817 0.2682 0.4454 0.2526 0.4206
GAME 0.3356 0.5734 0.3497 0.5938 0.3611 0.5636 0.3425 0.5384
MaskCLIP 0.2848 0.4885 0.2886 0.4957 0.3335 0.5351 0.3275 0.5267
CLIPSurgery 0.3115 0.5235 0.3217 0.5412 0.3832 0.6021 0.3727 0.5719
M2IB 0.3630 0.5953 0.3633 0.5951 0.3351 0.5411 0.3347 0.5410
Ours w/o $\lambda_{i}$ 0.2535 0.4379 0.2634 0.4568 0.3715 0.5831 0.3528 0.5556
Ours 0.2464 0.4272 0.2543 0.4420 0.3838 0.5993 0.3672 0.5749
Figure 8: Classification accuracy at Top-1 vs. (a) Deletion steps and (b) Insertion steps, on the ImageNet validation dataset with visual explanation heat maps from our Grad-ECLIP (solid) and other methods (dash).
TABLE II: Evaluation of image explanation faithfulness on MS COCO image-text retrieval (Karpathy’s split) validation dataset: AUC for Deletion and Insertion curves for performance on image retrieval (IR) and text retrieval (TR) tasks.
@@
IR TR IR TR
Method @1 @5 @1 @5 @1 @5 @1 @5
raw attention 0.1708 0.3554 0.1923 0.3720 0.1247 0.2552 0.1544 0.2969
Rollout 0.1948 0.3946 0.2268 0.4238 0.1294 0.2932 0.1753 0.3503
Grad-CAM 0.1717 0.3502 0.2161 0.4008 0.1027 0.2216 0.1152 0.2327
GAME 0.1706 0.3552 0.1982 0.3800 0.1537 0.3083 0.2097 0.3735
MaskCLIP 0.1321 0.2841 0.1516 0.2949 0.1423 0.2953 0.1891 0.3514
CLIPSurgery 0.1794 0.3652 0.2381 0.4292 0.1419 0.2941 0.1771 0.3384
M2IB 0.1797 0.3671 0.2057 0.3905 0.1469 0.3004 0.2058 0.3691
Ours w/o $\lambda_{i}$ 0.1390 0.2940 0.1827 0.3386 0.1403 0.2895 0.1735 0.3279
Ours 0.1246 0.2670 0.1550 0.2933 0.1576 0.3203 0.2056 0.3761
TABLE III: Evaluation of text explanation faithfulness on MS COCO image-text retrieval (Karpathy’s split) validation dataset: AUC for Deletion and Insertion curves with reporting image retrieval (IR) and text retrieval (TR) performance.
@@
Method IR TR IR TR
raw attention 0.2843 0.4917 0.0065 0.0328
Rollout 0.1221 0.2389 0.1052 0.2070
GAME 0.1083 0.2084 0.1146 0.2301
M2IB 0.2139 0.4256 0.0063 0.0375
Ours w/o $\lambda_{i}$ 0.1116 0.2113 0.1123 0.2361
Ours 0.0996 0.1770 0.1292 0.2536
4.2 Quantitative evaluation
In this section, we perform quantitative evaluations of Grad-ECLIP comparing with baselines. In §4.2.1, the explanation faithfulness is evaluated by the Deletion and Insertion metrics [75, 34, 36, 37, 43], which are also called perturbation tests [15, 16]. Moreover, in §4.2.2, we evaluate localization ability, when considering each visualization as a soft-segmentation of the image, using PointGame [76, 77] and segmentation tests [15]. Finally, in §4.2.3, we evaluate the processing time of Grad-ECLIP compared with other visual explanation methods. In this section, in order to understand the effect of the spatial importance term in (19), we also present Grad-ECLIP without using the spatial weight $\lambda_{i}$, i.e., setting $\lambda_{i}=1$, which is denoted as “w/o $\lambda_{i}$”.
4.2.1 Deletion and Insertion
A faithful explanation method should produce heat maps highlighting the important content in the image that has greatest impact on the model prediction. Deletion (negative perturbation) replaces input image pixels by random values step-by-step with the important pixels removed first based on the ordering of the heat map values, while recording the drop in prediction performance. Insertion adds image pixels to an empty image step-by-step based on the heat map importance, and records the performance increase. For deletion, larger drops are better, while for insertion larger increase is better. We consider each step as 0.5% of number of image pixels, and record results for $100$ steps. The model performance is measured using top-1 or top-5 zero-shot classification accuracy on the validation set of ImageNet [66] (ILSVRC) 2012, consisting of 50K images from 1000 classes. In particular, the perturbed image and each of the class names is fed into CLIP, and then classes with the highest scores are selected.
The insertion/deletion curves for top-1 accuracy are presented in Fig. 8, and the corresponding area under the curve (AUC) with top-1 and top-5 accuracy are presented in Tab. I. Steeper drop of performance with deletion steps corresponds to a lower deletion AUC, while quicker increase of performance with insert steps outputs a higher insertion AUC. Our method obtains the fastest performance drop for Deletion and largest performance increase for Insertion compared with most related works, showing that regions highlighted in our heat maps better represent explanations of CLIP. CLIPSurgery has comparable results to ours for Insertion, while performs poorly when evaluated with Deletion. The reason is that CLIPSurgery exhibits heat maps with nearly the same high values on the explained target region, so that the deletion operation fails to delete the most important pixels on the image at the beginning steps, which causes the deletion curve to decrease gradually, producing the high deletion AUC. Since CLIPSurgery can locate the explanation target with high values on the heat map, it performs well in the Insertion test. Our method without using the spatial importance (w/o $\lambda_{i}$) has slightly worse performance, but is still better than other baselines. As with [15, 16], we also use both the ground-truth class and the predicted class as the text prompt to generate heat maps, and our method is consistent with them, showing gains when using ground-truth text prompts.
@@ 4.2.1 Deletion and Insertion
We next evaluate the Deletion and Insertion performance for image and text retrieval tasks on the Karpathy’s validation split of MS COCO. To evaluate the image explanations, image pixels are removed or added step-by-step based on the text-specific explanation heat map. With the modified image replacing the original image, we record the image and text retrieval performance (recall @ top-1 and top-5 matching) to draw the deletion and insertion curve. Then, the AUC results of Deletion and Insertion with text-specific image explanations are reported in Tab. II. Grad-ECLIP surpasses the other methods on most metrics, which further demonstrates that our method produce high-quality visual explanation, regardless if the text is the class categories as in ImageNet or long captions as in MS COCO.
@@
Finally, we evaluate the faithfulness of our text explanations using the text version of Deletion and Insertion metric, where words are deleted or inserted based on the order of importance in the text heat map. Using images and caption annotations in MS COCO Karpathy’s split, we record the image-text retrieval performance for the modified caption, changing with total 5 steps with one word deleted/inserted at a time. The results in Tab. III show that Grad-ECLIP has the highest faithfulness (best deletion and insertion scores) compared with the other Transformer explanation methods. This demonstrates that Grad-ECLIP also has the excellent ability for image-specific text explanations.
4.2.2 Point Game and Segmentation Test
@@
We next evaluate the localization ability of the visual explanations. We adopt the ImageNet-Segmentation (ImageNet-S) [78] validation set, which has segment annotations on 12,419 images of 919 categories from ImageNet. Point Game (PG) is a commonly used metric to evaluate the localization correctness of a visual explanation. PG counts a hit score if the location with the largest value in the visual explanation heat map lies within the object region, which can be defined by the class segmentation mask. Then the PG accuracy is measured by averaging over all samples. Since PG only considers the maximum point, but not the spread of the heat map, we also conduct energy-PG [36], which calculates the proportion of heat map energy within the ground-truth mask versus the whole map. Similar to the evaluation by [15, 16], regarding the heat maps as soft-segmentation results, we adopt pixel accuracy (Pixel Acc.), average precision (AP), and averaged mask intersection over union (maskIoU) as additional metrics.
@@
The results for localization are shown in Tab. IV. Both versions of Grad-ECLIP significantly outperform other explanation methods on PG and energy-PG, which demonstrates that Grad-ECLIP can well reveal that the important pixels for CLIP are inside the object region. Comparing Grad-ECLIP with and without $\lambda_{i}$, Grad-ECLIP without $\lambda_{i}$ obtains relatively higher performance on pixel accuracy and maskIoU, since heat maps that contain more high-value pixels within the ground-truth mask have advantage on these two metrics. In Fig. 9(b,c), using $\lambda_{i}$ reduces the values on the mask while removing the surrounding noise. Due to the similar reason, CLIPSurgery obtains higher pixel accuracy and maskIoU, since it tends to put high heat map values on all the pixels of object region, and gets higher score when aggregating the heatmaps inside the object mask in these two evaluations. However, the lower PG, energy-PG and AP demonstrate that there are more high values generated outside of the object boundary. Better segmentation does not necessarily result in faithful explanations, in terms of both insertion and deletion metrics, as indicated in Table I.

### v2

@@ IV-A5 Applying Grad-ECLIP to explain CNN-based CLIP
Although the methodology in §III-B for Grad-ECLIP is derived from CLIP with Transformer architecture, here we show that our method is also applicable to CNN-based CLIP by using the attention layer in the final attention pooling. Figure 7 shows the visual explanation results for ResNet50-CLIP using Grad-ECLIP, and compared to other explanation methods that are compatible with CNN-based CLIP, including Grad-CAM, CLIPSurgery, and RISE. From the examples, our Grad-ECLIP is able to generate text-specific heat maps with less noise for interpreting CNN-based CLIP. In contrast to our method, Grad-CAM can produce similar highlights on the explained objects, but its heat map is noisier in the background. Similar to its performance on ViT-based CLIP, CLIPSurgery generates rougher heat maps, which tend to put high values on a coarse region of the object. Meanwhile, RISE fails to explain the CLIP model.
Figure 8: (left) An example showing the deletion and insertion process according to the Grad-ECLIP heat map on an image for 0 to 50 steps. The classification accuracy on each step is shown on the top, which forms the points of the deletion or insertion curves. (right) Classification accuracy at Top-1 vs. (a) Deletion steps and (b) Insertion steps, on the ImageNet validation dataset with visual explanation heat maps from our Grad-ECLIP (solid) and other methods (dashed).
IV-B Quantitative evaluation
In this section, we perform quantitative evaluations of Grad-ECLIP compared with baselines. In §IV-B1, the explanation faithfulness is evaluated by the Deletion and Insertion metrics [44, 8, 54, 53, 37], which are also called perturbation tests [10, 9]. Moreover, in §IV-B2, we evaluate the localization ability of CLIP using our Grad-ECLIP and the compared methods, when considering each visualization as a soft-segmentation of the image, using PointGame [69, 70] and segmentation tests [10]. Then, we conduct a user trust study to evaluate the trustability of the visual explanations and report the results in §IV-B3. Finally, in §IV-B4, we evaluate the processing time of Grad-ECLIP compared with other visual explanation methods. In this section, in order to understand the effect of the spatial importance term in (19), we also present Grad-ECLIP without using the spatial weight $\lambda_{i}$, i.e., setting $\lambda_{i}=1$, which is denoted as “w/o $\lambda_{i}$”.
IV-B1 Deletion and Insertion
The insertion/deletion process for an example image over the first 50 steps is displayed in Fig. 8 left. To summarize the insertion and deletion curves, we compute the area under the curve (AUC),
$\mathrm{Deletion}=\frac{1}{N_{step}}\sum_{n=1}^{N_{step}}Acc_{del}^{(n)},$ (21)
$\quad\mathrm{Insertion}=\frac{1}{N_{step}}\sum_{n=1}^{N_{step}}Acc_{ins}^{(n)},$ (22)
where $Acc_{ins}^{(n)}$ and $Acc_{del}^{(n)}$ are the model performance (e.g., accuracy) with inserted and deleted images at step $n$. For the deletion curve, a more faithful model will have larger drops in model performance (e.g., accuracy) as pixels are removed, and thus lower AUC for deletion curves is better. On the opposite, for the insertion curve, a more faithful model will have a larger increase in model performance as pixels are added, and thus a higher AUC for insertion curves is better. We also define an overall metric IMD (Insertion Minus Deletion), which summarizes the two AUCs:
$\mathrm{IMD}=\dfrac{1}{N_{step}}\sum_{n=1}^{N_{step}}Acc_{ins}^{(n)}-Acc_{del}^{(n)}.$ (23)
Higher IMD indicates better faithfulness, corresponding to higher insertion AUC and lower deletion AUC. We consider each step as 0.5% of the number of image pixels, and record results for $100$ steps. The model performance is measured using top-1 or top-5 zero-shot classification accuracy on the validation set of ImageNet [43] (ILSVRC) 2012, consisting of 50K images from 1000 classes. In particular, the perturbed image and each of the class names is fed into CLIP, and then classes with the highest scores are selected.
The insertion/deletion curves for top-1 accuracy are presented in Fig. 8 right, and AUC values for top-1 and top-5 accuracy are presented in Tab. I. Our method obtains the fastest performance drop for Deletion and the largest performance increase for Insertion compared with most related works (p<<.001), showing that regions highlighted in our heat maps better represent explanations of CLIP. CLIPSurgery has comparable results to ours for Insertion (p==0.374, 0.413, 0.074, 0.425), while performing poorly when evaluated with Deletion. The reason is that CLIPSurgery exhibits heat maps with nearly the same high values on the explained target region, so that the deletion operation fails to delete the most important pixels on the image at the beginning steps, which causes the deletion curve to decrease gradually, producing the high deletion AUC. Since CLIPSurgery can locate the explanation target with high values on the heat map, it performs well in the Insertion test. Overall, combining the deletion and insertion, our method has achieved the highest results in terms of IMD (p<<.001), which shows that changing image inputs based on Grad-ECLIP maps produces the greatest impact on the model prediction. Our method without using the spatial importance (w/o $\lambda_{i}$) has slightly worse performance, but is still better than other baselines. As with [10, 9], we also use both the ground-truth class and the predicted class as the text prompt to generate heat maps, and our method is consistent with them, showing gains when using ground-truth text prompts.
@@ IV-B1 Deletion and Insertion
TABLE I: Faithfulness evaluation of image explanation on the ImageNet validation set: Deletion, Insertion and IMD (Insertion minus Deletion), based on Top-1 (@1) or Top-5 (@5) classification accuracy. Either the ground-truth or the prediction is used as the text input into CLIP. The arrows indicate lower/higher value is better. The last row shows the p-value for a paired sample t-test between the top-2 methods (bolded and underlined), excluding Ours w/o $\lambda_{i}$. Significant values (p<<.05) are highlighted in blue.
@@
Ground-truth Prediction Ground-truth Prediction Ground-truth Prediction
Method @1 @5 @1 @5 @1 @5 @1 @5 @1 @5 @1 @5
raw attention 0.383 0.624 - - 0.249 0.419 - - -0.134 -0.204 - -
Rollout 0.408 0.656 - - 0.280 0.467 - - -0.128 -0.189 - -
Grad-CAM 0.342 0.563 0.352 0.582 0.268 0.445 0.253 0.421 -0.074 -0.117 -0.099 -0.161
GAME 0.336 0.573 0.350 0.594 0.361 0.564 0.343 0.538 0.025 -0.010 -0.007 -0.055
MaskCLIP 0.285 0.489 0.289 0.496 0.334 0.535 0.328 0.527 0.049 0.047 0.039 0.031
CLIPSurgery 0.312 0.524 0.322 0.541 0.383 0.602 0.373 0.572 0.070 0.079 0.051 0.030
M2IB 0.363 0.595 0.363 0.595 0.335 0.541 0.335 0.541 -0.028 -0.054 -0.029 -0.054
Ours 0.246 0.427 0.254 0.442 0.384 0.599 0.367 0.575 0.137 0.172 0.113 0.133
    - w/o $\lambda_{i}$ 0.254 0.438 0.263 0.457 0.372 0.583 0.353 0.556 0.118 0.145 0.089 0.099
p-value <<.001 <<.001 <<.001 <<.001 0.374 0.413 0.074 0.425 <<.001 <<.001 <<.001 <<.001
TABLE II: Evaluation of image explanation faithfulness on MS COCO image-text retrieval (Karpathy’s split) validation dataset: Deletion, Insertion and IMD with performance on image retrieval (IR) and text retrieval (TR) tasks. The last row shows the p-value for a paired sample t-test between the top-2 methods (bolded and underlined), excluding Ours w/o $\lambda_{i}$. Significant values (p<<.05) are highlighted in blue.
@@
IR TR IR TR IR TR
Method @1 @5 @1 @5 @1 @5 @1 @5 @1 @5 @1 @5
raw attention 0.171 0.355 0.192 0.372 0.125 0.255 0.154 0.297 -0.046 -0.038 -0.100 -0.075
Rollout 0.195 0.395 0.227 0.424 0.129 0.293 0.175 0.350 -0.065 -0.052 -0.128 -0.105
Grad-CAM 0.172 0.350 0.216 0.401 0.103 0.222 0.115 0.233 -0.069 -0.101 -0.129 -0.168
GAME 0.171 0.355 0.198 0.380 0.154 0.308 0.210 0.374 -0.017 0.012 -0.047 -0.006
MaskCLIP 0.132 0.284 0.152 0.295 0.142 0.295 0.189 0.351 0.010 0.038 0.011 0.057
CLIPSurgery 0.179 0.365 0.238 0.429 0.142 0.294 0.177 0.338 -0.037 -0.061 -0.071 -0.091
M2IB 0.179 0.367 0.206 0.391 0.147 0.300 0.206 0.369 -0.033 0.000 -0.067 -0.021
Ours 0.125 0.267 0.155 0.293 0.158 0.320 0.206 0.376 0.033 0.051 0.053 0.083
    - w/o $\lambda_{i}$ 0.139 0.294 0.183 0.339 0.140 0.289 0.174 0.328 0.001 -0.009 -0.005 -0.011
p-value <<.001 <<.001 0.106 0.705 0.011 0.001 0.297 0.645 <<.001 <<.001 0.001 0.001
TABLE III: Evaluation of text explanation faithfulness on MS COCO image-text retrieval (Karpathy’s split) validation dataset: Deletion, Insertion and IMD with performance on image retrieval (IR) and text retrieval (TR) tasks. The last row shows the p-value for a paired sample t-test between the top-2 methods (bolded and underlined), excluding Ours w/o $\lambda_{i}$. Significant values (p<<.05) are highlighted in blue.
@@
Method IR TR IR TR IR TR
raw attention 0.284 0.492 0.007 0.033 -0.278 -0.459
Rollout 0.122 0.239 0.105 0.207 -0.017 -0.032
GAME 0.108 0.208 0.115 0.230 0.006 0.022
M2IB 0.214 0.426 0.005 0.024 -0.233 -0.419
Ours 0.100 0.177 0.129 0.254 0.076 0.030
    - w/o $\lambda_{i}$ 0.112 0.211 0.112 0.236 0.001 0.025
p-value 0.019 0.017 0.025 0.026 0.019 0.019
@@
We next evaluate the Deletion and Insertion performance for image and text retrieval tasks on Karpathy’s validation split of MS COCO. To evaluate the image explanations, image pixels are removed or added step-by-step based on the text-specific explanation heat map. With the modified image replacing the original image, we record the image and text retrieval performance (recall @ top-1 and top-5 matching) to draw the deletion and insertion curve. Then, the AUC results of Deletion and Insertion with text-specific image explanations are reported in Tab. II. Grad-ECLIP surpasses the other methods on most metrics (p<<.001), which further demonstrates that our method produces high-quality visual explanation, regardless of whether the text is the class categories as in ImageNet or long captions as in MS COCO.
@@
Finally, we evaluate the faithfulness of our text explanations using the text version of the Deletion and Insertion metric, where words are deleted or inserted based on the order of importance in the text heat map. Using images and caption annotations in MS COCO Karpathy’s split, we record the image-text retrieval performance for the modified caption, changing with a total of 5 steps, with one word deleted/inserted at a time. The results in Tab. III show that Grad-ECLIP has the highest faithfulness (best deletion, insertion and IMD scores with p==0.019) compared with the other Transformer explanation methods.
The last row of Tables I, II, and III displays the paired sample t-test results conducted between the AUCs of the top-2 methods in the column, where pairs correspond to the same step on the two curves.
IV-B2 Point Game and Segmentation Test
Following the pointing game and segmentation test metric adopted in related works [54, 36, 57, 21, 69, 10], we evaluate the localization ability of CLIP using our Grad-ECLIP and the compared methods. It should be noted that these localization tests are mainly a rationalization of the CLIP prediction, rather than a faithfulness metric. Indeed, localization or segmentation performance is not necessarily indicative of the faithfulness, in terms of the insertion and deletion metrics, as indicated in Table I.
@@ IV-B2 Point Game and Segmentation Test
We adopt the ImageNet-Segmentation (ImageNet-S) [14] validation set, which has segment annotations on 12,419 images of 919 categories from ImageNet. Point Game (PG) is a commonly used metric to evaluate the localization correctness of a visual explanation. PG counts a hit score if the location with the largest value in the visual explanation heat map lies within the object region, which can be defined by the class segmentation mask. Then the PG accuracy is measured by averaging over all samples. Since PG only considers the maximum point, but not the spread of the heat map, we also conduct energy-PG [54], which calculates the proportion of heat map energy within the ground-truth mask versus the whole map. Similar to the evaluation by [10, 9], regarding the heat maps as soft-segmentation results, we adopt pixel accuracy (Pixel Acc.), average precision (AP), and averaged mask intersection over union (maskIoU) as additional metrics.
@@
The results for localization are shown in Tab. IV. Both versions of Grad-ECLIP significantly outperform other explanation methods on PG and energy-PG ($\chi^{2}$ test on PG hit numbers between Grad-ECLIP and the second best method CLIPSurgery, $\chi^{2}$=593.3, df=1, p<<.001), which demonstrates that Grad-ECLIP can well reveal that the important pixels for CLIP are inside the object region. Comparing Grad-ECLIP with and without $\lambda_{i}$, Grad-ECLIP without $\lambda_{i}$ obtains relatively higher performance on pixel accuracy and maskIoU, since heat maps that contain more high-value pixels within the ground-truth mask have an advantage on these two metrics. In Fig. 9(b,c) of the main paper, using $\lambda_{i}$ reduces the values on the mask while removing the surrounding noise. Due to a similar reason, CLIPSurgery obtains higher pixel accuracy and maskIoU, since it tends to put high heat map values on all the pixels of the object region, and gets a higher score when aggregating the heatmaps inside the object mask in these two evaluations. However, the lower PG, energy-PG, and AP demonstrate that there are more high values generated outside of the object boundary.
@@
    - w/o $\lambda_{i}$ 0.8356 0.4409 0.7365 0.5163 0.3314
IV-B3 User Study on Improving Understanding
In this section, we evaluate whether the Grad-ECLIP explanations can improve a user’s understanding of CLIP’s predictions. We adopt a forward simulation task [17], where the user is given an image and two captions and is asked to judge which caption will receive a higher matching score from the model under two conditions: with or without corresponding visual explanations provided. If the users’ judgment accuracy improves when provided with the explanations, then it suggests that the explanations allow the users’ to better understand the model’s predictions.
Ten images and their corresponding captions were randomly selected from the dataset. A questionnaire was created, where each image was displayed with two text caption options, and the participant was asked to select the caption that they think the CLIP model will give a higher matching score. The images were grouped into two sets of five, denoted as Set I and Set II. In Set I, the image and captions were displayed without explanations, and in Set II they were displayed with the corresponding Grad-ECLIP explanations for each image-caption pair. Before each Set, a few examples of image-caption pairs, their CLIP scores, and explanations (Set II only) were presented to the participant in a training phase. Set I was presented before Set II so that the participant’s initial understanding of the model was not biased by the explanations. The assignment of images to Set I and Set II was counterbalanced so that a given image appeared the same number of times in Set I and Set II over all participants. We collected responses from a total of 34 participants. More details about the questionnaires, including sample selection and examples, are presented in the Appendix.
The results of the user study indicate that the participants’ correct judgment of CLIP’s higher-matched sentence significantly improved when the Grad-ECLIP explanations were provided, from 67.6% to 80.0% accuracy (paired t-test, t(33)=1.904, p=0.032). Thus, the user study demonstrates that the Grad-ECLIP explanations can effectively assist users in understanding the prediction-making mechanism of the CLIP model, highlighting its superior image-text matching interpretability.